### v1

# DoFlow: Causal Generative Flows for Interventional and Counterfactual Time-Series Prediction

Abstract
Time-series forecasting increasingly demands not only accurate observational predictions but also causal forecasting under interventional and counterfactual queries in multivariate systems. We present DoFlow, a flow-based generative model defined over a causal DAG that delivers coherent observational and interventional predictions, as well as counterfactuals through the natural encoding–decoding mechanism of continuous normalizing flows (CNFs). We also provide a supporting counterfactual recovery result under certain assumptions. Beyond forecasting, DoFlow provides explicit likelihoods of future trajectories, enabling principled anomaly detection. Experiments on synthetic datasets with various causal DAG and real-world hydropower and cancer-treatment time series show that DoFlow achieves accurate system-wide observational forecasting, enables causal forecasting over interventional and counterfactual queries, and effectively detects anomalies. This work contributes to the broader goal of unifying causal reasoning and generative modeling for complex dynamical systems.
1 Introduction
Forecasting time-evolving systems is a foundational task in statistics and machine learning, with decades of progress across classical and modern paradigms. However, in many real-world systems, the goal extends beyond forecasting future observations: we also seek to answer interventional questions—“How will the forecast change under a planned action or modification of certain variables?”—and counterfactual queries—“What would the system have looked like had we intervened differently?” Yet most modern forecasting methods, despite their empirical success, are purely observational: they capture correlations but cannot reason about interventions or alternate scenarios. Classical causal inference methods for time series can model interventions, but they are often restricted to discrete treatments or low-dimensional systems and cannot generate full system trajectories. Bridging these regimes calls for a model that is both causally structured and generative, capable of generating probabilistically consistent trajectories under hypothetical actions.
In this paper, we address this gap by proposing DoFlow, a generative model based on Continuous Normalizing Flows (CNFs) that explicitly embeds a causal directed acyclic graph (DAG) structure. Leveraging the invertibility of CNFs and the temporal conditioning in Neural ODEs, DoFlow provides a unified framework for observational, interventional, and counterfactual forecasting. Beyond forecasting, DoFlow also yields explicit likelihoods for future trajectories, enabling principled change-point and anomaly detection. This framework unifies causal reasoning with generative modeling, advancing toward trustworthy inference and decision support for complex dynamical systems.
1.1 Related Work
To situate our work, we briefly review related lines of research spanning time-series forecasting and causal generative modeling.
  • Time-series forecasting. Modern approaches can be broadly categorized into four families: (i) classical statistical models such as ARIMA, state-space models, and VAR [69, 23, 26, 73, 59, 63]; (ii) deep sequence models including RNNs/LSTMs and attention-augmented variants [44, 22, 51, 38]; (iii) Transformer-based architectures offering scalability on long sequences [64, 72, 66, 48]; and (iv) recent flow and diffusion models that directly generate trajectories, including autoregressive [53, 54, 20], graph-augmented [14], and horizon-wide designs [1, 60]. These models typically condition on learned context representations and use expressive generative backbones to produce future forecasts, but they remain observational, lacking the ability to simulate interventional and counterfactual trajectories.
  • Causal generative modeling. For static data, interventional and counterfactual queries are traditionally addressed via structural equation models and do-calculus. Recent generative extensions include graph causal encoders [58, 57, 52], diffusion-based causal models [9], and model-agnostic counterfactual generators [30, 16]. However, these methods largely target non-temporal settings and do not capture causal dependencies that evolve over time. By contrast, DoFlow extends causal generative modeling to time-indexed DAGs, enabling coherent interventional and counterfactual forecasting for dynamic systems.
We next discuss three complementary research areas on causal time series—causal effect estimation, counterfactual explainability, and causal discovery—that further contextualize our contribution.
  • Causal effects on time series. Within a causal perspective, prior work has largely focused on modeling treatment effects and unobserved potential outcomes in sequential data, using methods such as Gaussian processes, variational Bayesian models, and classical time-series approaches like ARIMA and Interrupted Time Series [7, 59, 39, 4, 6, 5, 46, 21, 55, 27]. These studies typically examine how binary or discrete, fixed-time actions influence outcomes across a temporal horizon, often by extrapolating an alternative path and contrasting it with the observed post-intervention series. For time-varying treatments, variational approaches estimate individualized effects through balanced representations [36, 42, 34, 67]. Overall, treatment effects are often formalized as $\tau_{t}=\mathbb{E}[Y_{t}|A_{t-1}=j]-\mathbb{E}[Y_{t}|A_{t-1}=k]$, where $Y_{t}$ denotes the outcome and $A_{t}$ the time-varying action.
  • Counterfactual explainability. Beyond treatment effects, another line of work develops counterfactual explanations aimed at interpretability—identifying how past signals must be perturbed to alter a model’s predicted future trajectory or label. For example, “what adjustments to a patient’s breathing signal would lead the model to forecast deeper sleep stages?” Approaches include optimization-based perturbations [3, 17, 25], gradient- or motif-guided objectives [33, 65, 32], and self-interpretable models that couple prediction with counterfactual reasoning [68, 50]. While effective for explaining black-box predictors, these methods do not simulate system-wide trajectories under causal queries.
  • Causal discovery. Inferring causal directed acyclic graphs (DAGs) from observational time series has also been actively explored. Existing approaches include statistical dependence–based tests [45, 8], optimization-based formulations [49], and deep learning–based frameworks [62, 71, 12, 13], which collectively advance the recovery of temporal causal structures under various assumptions.
    Complementary to these efforts, there remains a need for models that explicitly embed causal structure within generative time-series forecasting, thereby enabling interventions on individual variables at arbitrary times and yielding coherent interventional and counterfactual predictions of system-wide trajectories. This direction—though relatively underexplored—is broadly important across domains. In hydropower systems [31], operators may ask how adjusting turbine flow over a given interval will propagate to downstream signals (interventional), or whether an alternative operating schedule could have avoided downstream anomalies (counterfactual). In healthcare [11, 28], dependencies among medication dosage, respiration, heart rate, and oxygenation motivate vital questions such as whether modified dosages during a specific window would have prevented abnormal oxygen levels over time.
2 Time-Conditioned Flow on DAG
2.1 Settings and goals
We consider a multivariate time series evolving over a causal directed acyclic graph (DAG) with nodes $\{1,\dots,K\}$ in a topologically sorted order. Each node $i$ in the DAG represents a univariate variable $X_{i,t}\in\mathbb{R}$ at discrete time step $t$. Let $\mathbf{X}_{t}:=\{X_{1,t},\dots,X_{K,t}\}$ denote the collection of all nodes at time $t$, and let $X_{\mathrm{pa}(i),t}:=\{X_{j,t}:j\in\mathrm{pa}(i)\}$ denote the values of node $i$’s parents at time $t$. The parent set $\mathrm{pa}(i)$ encodes the structural dependencies that remain fixed across time—for instance, in a physical system, upstream components serve as permanent causal parents of downstream ones. At each step $t$, the value of node $i$ is influenced by its own past trajectory $X_{i,t-}$ and the past trajectories of its parents $X_{\mathrm{pa}(i),t-}$:
$X_{i,t-}:=\{X_{i,s},s<t\},\quad X_{\mathrm{pa}(i),t-}:=\{X_{j,s},j\in\mathrm{pa}(i),s<t\}.$
Formally, we define the underlying structural causal model (SCM) as
$X_{i,t}:=f_{i}(X_{i,t-},X_{\mathrm{pa}(i),t-},U_{i,t}),$ (1)
where $U_{i,t}$ is an exogenous noise variable independent across nodes and time. Here, the SCM means that $X_{i,t}$ is generated by a function of its parental state $(X_{i,,t-},X_{\mathrm{pa}(i),t-})$ and an unobserved exogenous variable $U_{i,t}$. For training and evaluation, each sequence $\{\mathbf{X}_{1},\dots,\mathbf{X}_{T}\}$ is divided into a context window $\{\mathbf{X}_{1},\dots,\mathbf{X}_{\tau}\}$ used for conditioning, and a forecasting window $\{\mathbf{X}_{\tau+1},\dots,\mathbf{X}_{T}\}$ for prediction.
Goals.
Our goal is to develop time-series causal prediction under the following two types of regimes:
  • Interventional forecasting. We denote the intervention schedule by $\mathcal{I}\subseteq[K]\times\{\tau+1,\dots,T\}$, where each $(i,t)\in\mathcal{I}$ specifies an intervention on node $i$ at time $t$. Each intervention replaces the system’s natural evolution with a fixed value $\mathrm{do}(X_{i,t}:=\gamma_{i,t})$, collectively written as $\mathrm{do}(X_{\mathcal{I}}:=\gamma_{\mathcal{I}})$. This defines the interventional distribution:
    $p(\mathbf{X}_{\tau+1:T}\mid\mathbf{x}_{1:\tau},\,\mathrm{do}(X_{\mathcal{I}}:=\gamma_{\mathcal{I}})).$ (2)
    When $\mathcal{I}=\emptyset$, the model reduces to standard observational forecasting.
  • Counterfactual forecasting. Given a factual trajectory $\mathbf{x}_{\tau+1:T}^{\mathrm{F}}$, we aim to answer: “What would this specific sequence have looked like if variables $X_{\mathcal{I}}$ had instead been set to $\gamma_{\mathcal{I}}$?” This induces the counterfactual distribution:
    $p(\mathbf{X}_{\tau+1:T}^{\mathrm{CF}}\mid\mathbf{x}_{1:\tau},\,\mathbf{x}_{\tau+1:T}^{\mathrm{F}},\,\mathrm{do}(X_{\mathcal{I}}:=\gamma_{\mathcal{I}})).$ (3)
2.2 Continuous normalizing flow foundations
We propose DoFlow, a time-conditioned continuous normalizing flow (CNF) that autoregressively predicts the future value of each node. At each prediction step $t$, the CNF is conditioned on the past hidden states $H_{t-1}$, which encodes the temporal information up to time t1t{-}1. At each time-series step $t$, a CNF defines a continuous transformation between a base distribution (at s=0s{=}0, typically $\mathcal{N}(0,1)$) and the target distribution (at s=1s{=}1, corresponding to the data at time-series step $t$) through a neural ordinary differential equation (ODE):
$\small\frac{dx_{t}(s)}{ds}=v\!\left(x_{t}(s),s;\,H_{t-1}\right),\quad s\!\in\![0,1],$ (4)
where we learn the velocity field $v$ using a neural network. To predict $x_{t}$, one samples from the base distribution $z=x_{t}(0)\!\sim\!\mathcal{N}(0,1)$ and integrates the above ODE from $s=0$ to $s=1$. We train the velocity field $v$ using Conditional Flow Matching (CFM) loss [40], which directly regresses $v$ onto an analytically defined reference velocity field. Specifically, a reference path $\phi$ interpolates between a base sample $z$ and a data sample $x_{t}$ (typically using linear interpolation), whose derivative $\partial_{s}\phi$ defines the reference velocity. The model minimizes an $L_{2}$ loss between $v$ and $\partial_{s}\phi$, allowing efficient training while preserving the invertible mapping from base distribution to data distribution.
2.3 Time-conditioned continuous normalizing flow
RNN to summarize past histories.
We learn a continuous normalizing flow (CNF) for each node $i$ to autoregressively generate $X_{i,t}$ for $t=\tau+1,\dots,T$, conditioning each step on the node’s past history and its parents’ past histories. An RNN (can be either an LSTM or GRU [18]) is employed to summarize histories via a recurrent state $h_{i,t}$:
$h_{i,t}=\text{RNN}(\text{concat}\{x_{i,t},c_{i,t}\},h_{i,t-1}),$ (5)
where $c_{i,t}$ denotes optional conditioning variables for the time series, such as embeddings of temporal features. Therefore, for each node $i$ at time $t$, the flow is conditioned on the hidden states:
$H_{i,t-1}:=(h_{i,t-1},h_{\mathrm{pa}(i),t-1}),$ (6)
where $h_{i,t-1}$ and $h_{\mathrm{pa}(i),t-1}$ are the node and parents’ past recurrent states, respectively.
At forecasting time, we use $\hat{H}_{i,t}$ (for interventional forecasting) or $\hat{H}_{i,t}^{\mathrm{CF}}$ (for counterfactual forecasting) to denote the hidden states that are updated from the model predicted values. Besides, $H_{i,t}^{\mathrm{F}}$ denotes the factual hidden state computed from the observed factual trajectory.
Time-conditioned continuous normalizing flow.
Conditioned on histories $H_{i,t-1}$, we train a separate CNF for each node $i$ that are shared across time-series steps $t$. The Neural ODE of the conditional continuous normalizing flow is defined as:
$\small\frac{dx_{i,t}(s)}{ds}=v_{i}(x_{i,t}(s),s;H_{i,t-1}),\quad s\in[0,1],\quad t\in\{\tau+1,\tau+2,\dots,T\},$ (7)
which transforms samples from the base distribution $\mathcal{N}(0,1)$ to the time-series distribution at time $t$. To predict $x_{i,t}$, one samples from the base distribution $z=x_{i,t}(0)\!\sim\!\mathcal{N}(0,1)$ and integrates the ODE from s=0s{=}0 to s=1s{=}1.
Training loss.
At the start of the forecasting window, $H_{i,\tau}$ is initialized from the context sequence $X_{i,1:\tau}$. For later forecasting steps $t\geq\tau+1$, the hidden states $H_{i,t}$ are updated recursively during training using the observed values $(X_{i,t},X_{\mathrm{pa}(i),t})$. The training loss is calculated over the entire forecasting window. For each $t\in\{\tau+1,\dots,T\}$, we define the reference path $\phi$ as a straight-line interpolation between the training sample $x_{i,t}$ and Gaussian noise $z\sim\mathcal{N}(0,I)$:
$\phi(x_{i,t},z;s):=(1-s)\,x_{i,t}+s\,z,\quad\partial_{s}\phi(x_{i,t},z;s)=z-x_{i,t},$ (8)
with $s\sim\mathcal{U}[0,1]$.
Here, $\partial_{s}\phi$ represents the reference velocity field, which the model’s learned velocity field $v_{i}$ aims to approximate. Training thus minimizes the squared $L_{2}$ distance between $v_{i}$ and $\partial_{s}\phi$. Therefore, the training loss of the flow using conditional flow matching [40] becomes:
$\small\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{\mathbf{X}_{1:T}\sim p_{\mathcal{X}}}\Bigg[\frac{1}{K(T-\tau)}\sum_{i=1}^{K}\sum_{t=\tau+1}^{T}\;\mathbb{E}_{s\sim\mathcal{U}[0,1],\,z\sim\mathcal{N}(0,I)}\big\|v_{i}\big(\phi(x_{i,t},z;s),\,s;\,H_{i,t-1}\big)-\partial_{s}\phi(x_{i,t},z;s)\big\|_{2}^{2}\Bigg],$ (9)
where $\theta$ encompasses both the parameters of the velocity field $v_{i}$ and the RNN parameter.
After training, we obtain a velocity field for each node $i$ on the causal DAG. We next define the forward and reverse processes of the Neural ODE.
Forward process.
We treat the forward process as an encoding operation, denoted by the function $Z_{t}:=\Phi_{\theta}(X_{t};\,H_{t})$. Given the velocity field $v_{i}$, the forward process pushes an observed factual outcome $x_{i,t}^{\mathrm{F}}$ (at $s=0$) to a latent embedding $z_{i,t}^{\mathrm{F}}$ (at $s=1$), conditioned on a factual hidden state $H_{i,t-1}^{\mathrm{F}}$ that summarizes the past observed factual data. Formally,
$z_{i,t}^{\mathrm{F}}=\Phi_{\theta}(x_{i,t}^{\mathrm{F}};\,H_{i,t}^{\mathrm{F}})=x_{i,t}^{\mathrm{F}}+\int_{0}^{1}\!v_{i}\big(x_{i,t}(s),\,s;\,H_{i,t-1}^{\mathrm{F}}\big)\,ds,\ \text{with}\ x_{i,t}(0)=x_{i,t}^{\mathrm{F}}.$ (10)
Through this process, $z_{i,t}^{\mathrm{F}}$ serves as a latent encoding of the factual observation $x_{i,t}^{\mathrm{F}}$.
Reverse process.
We treat the reverse process as a decoding operation. It is initialized differently for interventional (Section 3.1) and counterfactual prediction (Section 3.2). In general, to predict the value of node $i$ at time $t$, given a latent representation $z_{i,t}$, which can be either sampled from $\mathcal{N}(0,1)$ or obtained by encoding a factual sample, the reverse process is defined as:
$\hat{x}_{i,t}=\Phi_{\theta}^{-1}\!\big(z_{i,t};\,\hat{H}_{i,t-1}\big)=z_{i,t}-\!\!\int_{0}^{1}\!v_{i}\big(x_{i,t}(s),\,s;\,\hat{H}_{i,t-1}\big)\,ds,\ \text{with}\ x_{i,t}(1)=z_{i,t}.$ (11)
Here, $\hat{H}_{i,t}$ is autoregressively updated using the predicted values $\hat{x}_{i,t}$ and $\hat{x}_{\mathrm{pa}(i),t}$, and serves as the conditioning state for the next time step.
3 Interventional and Counterfactual Predictions
Figure 1: (A) RNN State Update. (B) Observational/Interventional Forecasting. Forecasts are generated by decoding from latent $z_{i,t}\sim N(0,1)$, conditioned on $\hat{H}_{i,t-1}$ updated with the last predicted $(\hat{x}_{i,t-1},\hat{x}_{\mathrm{pa}(i),t-1})$. (C) A factual observation $x_{i,t}^{\mathrm{F}}$ is encoded with its factual state $H_{i,t}^{\mathrm{F}}$ into $z_{i,t}^{\mathrm{F}}$, then decoded under the counterfactual state $\hat{H}_{i,t-1}^{\mathrm{CF}}$ to yield $\hat{x}_{i,t}^{\mathrm{CF}}$. Factual states $H_{i,t-1}^{\mathrm{F}}$ are updated from observed $x_{i,t-1}^{\mathrm{F}}$, while counterfactual states $\hat{H}_{i,t-1}^{\mathrm{CF}}$ are updated from the previously generated $\hat{x}_{i,t-1}^{\mathrm{CF}}$.
3.1 Observational and interventional prediction
At inference time, we forecast each node’s value one step ahead using the reverse process conditioned on the latest hidden states. Assume that our intervention schedule is $\mathcal{I}$ with intervened values $\{\gamma_{i,t}\}$, and the purely observational case is given by $\mathcal{I}=\emptyset$.
The forecasting at time step $t$ is proceeded in a topologically sorted order, meaning that parent nodes are forecasted first, followed by their children. For a node $i$ intervened at time $t$ with value $\gamma_{i,t}$, the forecast is fixed to the intervened value, i.e., $\hat{x}_{i,t}\leftarrow\gamma_{i,t}$. For non-intervened nodes, the forecast $\hat{x}_{i,t}$ is generated by the flow model using reverse process (Eq. 11), conditioned on its hidden state $\hat{H}_{i,t-1}$, which is updated using the previous forecasts $\hat{x}_{i,t-1}$ and $\hat{x}_{\mathrm{pa}(i),t-1}$. The overall procedure for observational and interventional forecasting over the time series is summarized in Algorithm 1. An illustrative figure is shown in Panel B of Figure 1.
1:Input: Context window $\{x_{i,1:\tau}\}_{i=1}^{K}$; intervention schedule $\mathcal{I}$ with values $\{\gamma_{i,t}\}$
2:Initialize hidden states $\hat{H}_{i,\tau}$ with $x_{i,1:\tau}$ for all $i=1,\dots,K$
3:for $t=\tau+1$ to $T$ do
4:for $i=1,\dots,K$ do {topological order}
5:  if $(i,t)\in\mathcal{I}$ then
6:   $\hat{x}_{i,t}\leftarrow\gamma_{i,t}$
7:  else
8:   Sample $z_{i,t}\sim\mathcal{N}(0,1)$
9:   $\hat{x}_{i,t}\leftarrow\Phi_{\theta}^{-1}\!\big(z_{i,t};\,\hat{H}_{i,t-1}\big)$ {Eq. (11)}
10:  end if
11:  $h_{i,t},h_{\mathrm{pa}(i),t}\xleftarrow[\text{}]{\text{update}}(\hat{x}_{i,t},\;\hat{x}_{\mathrm{pa}(i),t})$
12:  $\hat{H}_{i,t}\leftarrow(h_{i,t},h_{\mathrm{pa}(i),t})$
13:end for
14:end for
15:Output: $\{\hat{x}_{i,t}\}_{i=1..K,\;t=\tau+1,..,T}$
Algorithm 1 Time Series Observational/Interventional Forecasting
3.2 Time series forecasting for counterfactual queries
Counterfactual forecasting follows the standard abduction–action–prediction procedure. Given a factual trajectory $\{x_{i,\tau+1:T}^{\mathrm{F}}\}_{i=1}^{K}$, we proceed as follows: (i) abduction – we infer latent variables by encoding each observed factual value into its latent representation $z_{i,t}^{\mathrm{F}}$ through the forward process (10), conditioned on a factual hidden state $H_{i,t-1}^{\mathrm{F}}$ that summarizes past factual observations; (ii) action – we apply the specified intervention schedule $\mathcal{I}$ with values $\{\gamma_{i,t}\}$ to the corresponding nodes; and (iii) prediction – we autoregressively generate the counterfactual trajectory $\{\hat{x}_{i,t}^{\mathrm{CF}}\}$ through the reverse process (11), starting from the abducted latent representations $\{z_{i,t}^{\mathrm{F}}\}$.
We first compute the factual hidden states $\{H^{\mathrm{F}}_{i,t}\}_{t=\tau}^{T-1}$ from the observed context $\{x_{i,1:\tau}\}$ and the observed factual trajectory $\{x^{\mathrm{F}}_{i,\tau+1:T}\}$. These factual states are used only for encoding factual values into their latent representations, which serve as the starting point for counterfactual decoding. At each time step $t$, nodes are predicted in topological order. For $(i,t)\in\mathcal{I}$, the counterfactual forecast is set as $\hat{x}_{i,t}^{\mathrm{CF}}:=\gamma_{i,t}$, and the counterfactual hidden state $\hat{H}_{i,t}^{\mathrm{CF}}$ is updated accordingly. For non-intervened nodes, the factual value $x_{i,t}^{\mathrm{F}}$ is first encoded as $z_{i,t}^{\mathrm{F}}=\Phi_{\theta}(x_{i,t}^{\mathrm{F}};\,H_{i,t-1}^{\mathrm{F}})$, then decoded under the counterfactual hidden state as $\hat{x}_{i,t}^{\mathrm{CF}}=\Phi_{\theta}^{-1}(z_{i,t}^{\mathrm{F}};\,\hat{H}_{i,t-1}^{\mathrm{CF}})$, followed by updating $\hat{H}_{i,t}^{\mathrm{CF}}$ with the generated value and its parents. The overall counterfactual generation procedure is summarized in Algorithm 2 and illustrated in Panel C of Figure 1.
1:Input: Context window $\{x_{i,1:\tau}\}_{i=1}^{K}$; factual sample $\{x_{i,\tau+1:T}^{\text{F}}\}_{i=1}^{K}$; intervention schedule $\mathcal{I}$ with values $\{\gamma_{i,t}\}$
2:Obtain factual hidden states $\{H^{\mathrm{F}}_{i,t}\}_{t=\tau}^{T-1}$ from context $\{x_{i,1:\tau}\}$ and observed factual $\{x^{\mathrm{F}}_{i,\tau+1:T}\}$
3:Initialize counterfactual hidden states $\hat{H}^{\mathrm{CF}}_{i,\tau}=H_{i,\tau}$ with context $\{x_{i,1:\tau}\}$ for all $i=1,\dots,K$
4:for $t=\tau+1$ to $T$ do
5:for $i=1,\dots,K$ do {nodes in topological order}
6:  if $(i,t)\in\mathcal{I}$ then
7:   $\hat{x}_{i,t}^{\text{CF}}\leftarrow\gamma_{i,t}$
8:  else
9:   $z_{i,t}^{\text{F}}\leftarrow\Phi_{\theta}(x_{i,t}^{\text{F}},H_{i,t-1}^{\text{F}})$   {Eq. (10); Abduction}
10:   $\hat{x}_{i,t}^{\text{CF}}\leftarrow\Phi_{\theta}^{-1}(z_{i,t}^{\text{F}},\hat{H}_{i,t-1}^{\text{CF}})$ {Eq. (11); Action-Prediction}
11:  end if
12:  $h_{i,t},h_{\mathrm{pa}(i),t}\xleftarrow[\text{}]{\text{update}}(\hat{x}_{i,t}^{\mathrm{CF}},\;\hat{x}_{\mathrm{pa}(i),t}^{\mathrm{CF}})$
13:  $\hat{H}_{i,t}^{\text{CF}}\leftarrow(h_{i,t},h_{\mathrm{pa}(i),t})$
14:end for
15:end for
16:Output: $\{\hat{x}_{i,t}^{\text{CF}}\}_{i=1..K,\;t=\tau+1,..T}$
Algorithm 2 Counterfactual Time Series Generation
3.3 Additional property: likelihood-based anomaly detection
Another advantage of our framework is its ability to learn the density of the future trajectory over the forecasting window. For a single node (index $i$ omitted for clarity), we denote its density by $p_{\theta,X_{\tau+1:T}}(\cdot|\hat{H}_{\tau},z_{\tau+1:T})$, which represents the density of the future outcomes learned by the model, conditioned on the context state $\hat{H}_{\tau}$ and a sequence of simulated base samples $z_{\tau+1:T}$. The explicit form of the learned density, mapping from latent $z_{\tau+1:T}$ to data $\hat{x}_{\tau+1:T}$, is given as:
Proposition 3.1.
Given base samples $z_{\tau+1:T}\sim q(\cdot)$, the log-density of the generated time series obtained via the continuous normalizing flow is:
$\small\log p_{\theta,X_{\tau+1:T}}\!\left(\hat{x}_{\tau+1:T}\mid\hat{H}_{\tau},z_{\tau+1:T}\right)=\sum_{t=\tau+1}^{T}\left[\log q\!\left(z_{t}\right)+\int_{0}^{1}\nabla\!\cdot\,v_{\theta}\bigl(x_{t}(s),s;\,\hat{H}_{t-1}\bigr)\,ds\right].$ (12)
Anomaly Detection.
Since anomalies deviate substantially from normal patterns, we expect that they have lower densities. Therefore, the log-density as computed in (12) naturally provides a principled measure of anomaly.
4 Theoretical Properties
In this section, we present a theoretical result on the counterfactual recovery of our algorithm under certain assumptions. For simplicity, we fix and omit the index $i$ from the notation. Recall that the encoding function is defined as $\Phi_{\theta}:\mathcal{X}\times\mathcal{H}\to\mathcal{Z}$ and the decoding function as $\Phi_{\theta}^{-1}:\mathcal{Z}\times\mathcal{H}\to\mathcal{X}$, corresponding to (10) and (11), respectively. Recall that the underlying structural causal model is given by $X_{t}:=f(X_{t-},X_{\text{pa},t-},U_{t})$. In the following, we present a supporting result on the counterfactual recovery properties of DoFlow. We begin by introducing the following assumptions.
Assumption 4.1.
(A1) $U_{t}\perp\!\!\!\perp(X_{t-},X_{\text{pa},t-})$.
(A2) The structural causal equation $f(\cdot,U_{t})$ is monotone in $U_{t}$.
(A3) For the encoded latent variable $Z_{t}=\Phi_{\theta}(X_{t};H_{t-1})$, the conditional distribution satisfies $p_{\theta}(Z_{t}\mid H_{t-1})=q(Z_{t})=N(Z_{t};0,1)$.
Remark 4.2.
Each node $X_{i}\in\mathbb{R}$, so (A2) is assumed in the univariate case. It is automatically satisfied under additive SCMs, i.e., $X_{t}=f^{*}(X_{t-},X_{\text{pa},t-})+U_{t}$. Under certain identifiability conditions, it also holds for non-linear models [70, 61]. Our theoretical results apply only to monotone SCM settings. One may notice that (A1)–(A2) mirror those in Bijective Generation Mechanisms (BGM) [47], which establish model-agnostic identifiability with an additional assumption on distribution matching. In contrast, our Corollary 4.5 provides model-specific, pointwise recovery for our CNF under (A1)–(A3) without requiring distribution matching; see Appendix D.
Note that (A3) implies that the encoded $Z_{t}=\Phi_{\theta}(X_{t};H_{t-1})$ is statistically independent of $H_{t-1}$, and equivalently of $(X_{t-},X_{\text{pa},t-})$, in distribution. In the infinite-data limit with exact training, the continuous normalizing flow maps every $X_{t}$, conditioned on any fixed $(X_{t-},X_{\text{pa},t-})$, to the same base distribution $q(Z_{t})=N(0,1)$, so that (A3) holds exactly. In finite-sample regimes, deviations may occur; we empirically assess the validity of (A3) in Appendix E.1.
Under Assumption 4.1, we present the first result in this paper:
Proposition 4.3 (Encoded as a function of the exogenous noise $U_{t}$).
Let Assumption 4.1 hold. Without loss of generality, suppose the exogenous noise $U_{t}\sim\mathrm{Unif}[0,1]$. At each time $t$, the observed variable is generated by the structural causal model $X_{t}=f(X_{t-},X_{\mathrm{pa},t-},U_{t})$, and that the flow encoder produces $Z_{t}=\Phi_{\theta}(X_{t};H_{t-1})$. Then there exists a continuously differentiable bijection $g:\mathcal{U}\to\mathcal{Z}$, functionally invariant to $H_{t-1}$, such that,
$Z_{t}\;=\;\Phi_{\theta}\bigl(X_{t};H_{t-1}\bigr)\;=\;\Phi_{\theta}\bigl(f(X_{t-},X_{\mathrm{pa},t-},U_{t});H_{t-1}\bigr)\;=\;g\bigl(U_{t}\bigr)\quad\text{a.s.}$ (13)
Remark 4.4.
Proposition 13 states that $Z_{t}$ is a function of the exogenous noise $U_{t}\sim\mathrm{Unif}[0,1]$. This assumption can be relaxed to other noise distributions. For example, if $Z\sim\mathcal{N}(0,1)$ with CDF $F$, then $U=F(Z)\sim\mathrm{Unif}[0,1]$, and any assignment $f(\cdot,U)$ can equivalently be written as $\tilde{f}(\cdot,Z)=f(\cdot,F(Z))$. Therefore, the theorem’s conclusions also hold under Gaussian noise.
Following Proposition 13, we state a counterfactual recovery result under monotone SCMs. Given an intervention schedule $\mathcal{I}$ with values $\{\gamma_{t}\}$, the true counterfactual process is defined recursively as:
$\small X^{\mathrm{CF}}_{t}\;=\;\begin{cases}\gamma_{t},&t\in\mathcal{I},\\[2.0pt] f\!\bigl(X^{\mathrm{CF}}_{t-},\,X^{\mathrm{CF}}_{\mathrm{pa},t-},\,U_{t}\bigr),&\text{otherwise,}\end{cases}$ (14)
where $U_{t}$ is abducted from the factual sample $X_{t}^{\mathrm{F}}$. In contrast, Algorithm 2 updates the estimated $\hat{H}_{t-1}^{\mathrm{CF}}$ recursively (Line 13), using the estimates $\hat{X}_{t-1}^{\mathrm{CF}}$ and $\hat{X}_{\mathrm{pa},t-1}^{\mathrm{CF}}$. We can now state the following result:
Corollary 4.5 (Counterfactual recovery).
Let Assumption 4.1 hold. Consider a factual sample generated by the structural causal model $X_{t}^{\mathrm{F}}=f(X_{t-},X_{\mathrm{pa},t-},U_{t})$, and let its encoded latent be $Z_{t}^{\mathrm{F}}:=\Phi_{\theta}\bigl(X_{t}^{\mathrm{F}};H_{t-1}^{\mathrm{F}}\bigr)$. At time step $t$, we apply the intervention $\mathrm{do}\bigl(X_{t-}=\hat{X}_{t-}^{\mathrm{CF}},X_{\mathrm{pa},t-}=\hat{X}_{\mathrm{pa},t-}^{\mathrm{CF}}\bigr)$, yielding the counterfactual hidden state $\hat{H}_{t-1}^{\mathrm{CF}}$. Then the decoder recovers the true counterfactual at time step $t$ almost surely:
$\hat{X}_{t}^{\mathrm{CF}}:=\Phi_{\theta}^{-1}\!\bigl(Z_{t}^{\mathrm{F}};\,\hat{H}_{t-1}^{\mathrm{CF}}\bigr)=X_{t}^{\mathrm{CF}}.$
5 Experiments
In this section, we evaluate DoFlow on observational forecasting, causal forecasting under both interventional and counterfactual queries, and anomaly detection, using both synthetic and real-world datasets.
Because in real-world settings, the ground-truth counterfactual is never observable, and the ground-truth interventional is only observable if interventions are actively conducted under a correct causal DAG, we rely on synthetic experiments to obtain quantitative performance metrics. For real-world evaluation, we assess our model on interventional queries and anomaly detection using hydropower datasets from Argonne National Laboratory, and on interventional treatment effect estimation using the cancer-treatment dataset from [5].
5.1 Synthetic Data Experiments
We consider both the additive noise model, i.e., $f_{i}(X_{i,t-},X_{\mathrm{pa}(i),t-},U_{i,t})=f_{i}^{*}(X_{i,t-},X_{\mathrm{pa}(i),t-})+U_{i,t},$ which satisfies Assumption 4.1 and supports the counterfactual recovery result, as well as more general non-linear and non-additive (NLNA) cases of $f_{i}(X_{i,t-},X_{\mathrm{pa}(i),t-},U_{i,t})$ to test the model’s robustness beyond the scope of our counterfactual recovery result.
We evaluate the model on multiple structurally diverse causal DAGs, using Root Mean Squared Error (RMSE) and Maximum Mean Discrepancy (MMD) as evaluation metrics. Since DoFlow generates a single counterfactual forecasting trajectory aimed at recovering the true counterfactual, MMD is not applicable. Therefore, we report only RMSE for counterfactual forecasting. Detailed simulation setups and metric definitions are provided in Appendix C.
For baseline comparisons in observational forecasting, we consider: a pure RNN-based method, the Gated Recurrent Unit (GRU) [10]; transformer-based methods, including the Temporal Fusion Transformer (TFT) [37] and the Time Series Dense Encoder (TiDE) [15]; and an all-MLP method with a specialized contextual mixing structure, the Time Series Mixer (TSMixer) [19]. We also compare against probabilistic methods, including DeepVAR [56] — a deep RNN-based model with multivariate Gaussian outputs, and MQF2 [29] — a convex deep neural network that learns multivariate quantile functions. We extend our acknowledgements to the Python packages Darts [24] and GluonTS [2], which we use to directly test several modern baselines in this paper.
To our knowledge, comparatively few works tackle forecasting interventional/counterfactual time-series on a causal DAG. Nevertheless, we can adapt strong observational forecasters for interventional simulations by training a separate model for each node in the causal DAG, conditioned on its parents. However, counterfactual generation remains challenging for these baselines.
Table 1: RMSE for observational, interventional, and counterfactual time series forecasting across causal structures: Tree, Diamond, and FC-Layer. Results for the Chain structure are provided in Table 6 (Appendix). MMD results are provided in Table 5 (Appendix). Reported values are averaged over 50 test batches, each containing 128 test series.
Tree Diamond FC-Layer
Additive NLNA Additive NLNA Additive NLNA
Obs. Int. CF. Obs. Int. CF. Obs. Int. CF. Obs. Int. CF. Obs. Int. CF. Obs. Int. CF.
DoFlow $\mathbf{0.57_{\pm.09}}$ $\mathbf{0.66_{\pm.11}}$ $\mathbf{0.65_{\pm.10}}$ $\mathbf{0.59_{\pm.13}}$ $\mathbf{0.68_{\pm.14}}$ $\mathbf{0.69_{\pm.14}}$ $0.52_{\pm.10}$ $\mathbf{0.71_{\pm.14}}$ $\mathbf{0.72_{\pm.14}}$ $\mathbf{0.31_{\pm.07}}$ $\mathbf{0.63_{\pm.14}}$ $\mathbf{0.69_{\pm.15}}$ $0.39_{\pm.12}$ $\mathbf{0.41_{\pm.13}}$ $\mathbf{0.49_{\pm.13}}$ $\mathbf{0.54_{\pm.09}}$ $\mathbf{0.73_{\pm.15}}$ $\mathbf{0.64_{\pm.12}}$
GRU $0.65_{\pm.08}$ $1.01_{\pm.10}$ NA $0.63_{\pm.07}$ $1.04_{\pm.11}$ NA $0.58_{\pm.06}$ $0.94_{\pm.11}$ NA $0.37_{\pm.05}$ $0.99_{\pm.12}$ NA $\mathbf{0.38_{\pm.05}}$ $0.72_{\pm.10}$ NA $0.58_{\pm.07}$ $1.05_{\pm.13}$ NA
TFT $\mathbf{0.58_{\pm.11}}$ $0.97_{\pm.17}$ NA $0.63_{\pm.15}$ $1.01_{\pm.18}$ NA $0.63_{\pm.17}$ $1.18_{\pm.21}$ NA $0.40_{\pm.08}$ $1.09_{\pm.20}$ NA $0.47_{\pm.14}$ $0.83_{\pm.16}$ NA $0.62_{\pm.15}$ $1.02_{\pm.23}$ NA
TiDE $0.60_{\pm.13}$ $1.15_{\pm.21}$ NA $0.68_{\pm.14}$ $1.13_{\pm.20}$ NA $\mathbf{0.50_{\pm.12}}$ $1.05_{\pm.19}$ NA $0.33_{\pm.10}$ $0.99_{\pm.16}$ NA $0.43_{\pm.12}$ $0.75_{\pm.14}$ NA $0.66_{\pm.17}$ $1.10_{\pm.20}$ NA
TSMixer $0.63_{\pm.13}$ $1.08_{\pm.18}$ NA $0.65_{\pm.13}$ $1.07_{\pm.18}$ NA $\mathbf{0.49_{\pm.10}}$ $1.12_{\pm.20}$ NA $0.35_{\pm.11}$ $0.97_{\pm.15}$ NA $0.42_{\pm.11}$ $0.79_{\pm.15}$ NA $0.61_{\pm.15}$ $1.13_{\pm.19}$ NA
DeepVAR $0.64_{\pm.07}$ $0.74_{\pm.12}$ NA $0.65_{\pm.09}$ $0.86_{\pm.15}$ NA $0.68_{\pm.09}$ $0.86_{\pm.17}$ NA $0.45_{\pm.08}$ $0.94_{\pm.16}$ NA $0.54_{\pm.10}$ $1.17_{\pm.18}$ NA $0.69_{\pm.11}$ $1.57_{\pm.21}$ NA
MQF2 $\mathbf{0.58_{\pm.10}}$ $1.23_{\pm.19}$ NA $0.67_{\pm.11}$ $1.30_{\pm.21}$ NA $0.64_{\pm.12}$ $1.20_{\pm.16}$ NA $0.38_{\pm.06}$ $1.17_{\pm.18}$ NA $0.50_{\pm.09}$ $1.09_{\pm.12}$ NA $0.57_{\pm.10}$ $1.33_{\pm.22}$ NA
Tables 1 and 5 (Appendix) report the RMSE and MMD results for observational, interventional, and counterfactual time-series forecasting across multiple causal structures (detailed in Appendix C). DoFlow consistently delivers strong performance in standard observational forecasting and interventional forecasting compared with our adapted baselines, and uniquely supports counterfactual forecasting. We present one visual result of interventional and counterfactual in Figure 2.
Figure 2: Left:Layer” interventional forecasting results. Nodes $X_{1,t}$, $X_{2,t}$, and $X_{3,t}$ are intervened. DoFlow provides 50% and 90% prediction intervals; the orange lines indicate the true interventional future. Right:Tree” counterfactual forecasting results. Node $X_{1,t}$ is intervened. DoFlow provides a single forecast in green; the orange lines indicate the true counterfactual future.
5.2 Real Application: Hydropower System
We evaluate DoFlow on real-world hydropower time-series data from Argonne National Laboratory. In this system, water drives a turbine that powers a generator and passes through a transformer before reaching the grid, with control systems monitoring the process. Signals such as water flow, vibration, and electric current are recorded from each component, forming a natural as shown in Fig. 11.
Our evaluation focuses on two tasks: (1) whether DoFlow can accurately forecast each component’s time series under interventional queries; and (2) whether DoFlow can accurately detect power outages in advance using log-density, as discussed in Section 3.3.
Figure 3 illustrates interventional forecasting in the hydropower system during a true power outage. At this point, the turbine signals ($X_{1,t},X_{2,t}$) break down, leading the entire system into a forced outage. Notably, for $X_{3,t}$ and $X_{4,t}$, DoFlow successfully predicts the characteristic “spikes” in the generator signals that follow the turbine failure, demonstrating that our DAG-based approach captures the complex causal relationships across system. Quantitative metrics, together with the anomaly detection visualizations from the second task, are provided in the Appendix E.4 due to page limits.
Figure 3: Hydropower – Interventional.
5.3 Real Application: Cancer Treatment Outcomes
Our DoFlow can naturally model interventional forecasting for causal treatment effect estimation in time series. We evaluate it on the cancer-treatment benchmark of Bica et al. [5], which contains daily patient-level tumor volumes and administered therapies. At each day $t$, the treatment variables $\mathbf{X}_{t}=\{X_{i,t}\}_{i=1}^{4}$ (chemotherapy assignment, radiotherapy assignment, chemotherapy dosage, and radiotherapy dosage) act as causal parents of the outcome $Y_{t}$ (tumor volume) on the causal DAG. For each test patient, we use the first $55$ days as observational context and perform interventional rollouts for the next $5$ days under ten distinct treatment options, yielding predicted outcomes $\hat{Y}_{t+1:t+5}$. The normalized RMSE between the true $Y_{t+1:t+5}$ and predicted $\hat{Y}_{t+1:t+5}$ is reported in Appendix E.5.
The detailed settings and results are provided in Appendix E.5. Notably, DoFlow achieves substantial improvements over prior baselines on this causal treatment effect task, as summarized in Table 8.
5.4 Results and Comparisons
  • Training Efficiency. DoFlow trains a separate flow per node in the causal DAG, but each network can be much shallower than transformer models. As shown in Tables 3 and 4 (Appendix), its total model size remains modest, and its training and sampling times are comparable to modern baselines.
  • Interventional Performance. We adapted modern baselines for conditional interventional generation, but DoFlow performs notably better: its RNN–flow design jointly encodes causal histories and propagates interventions through recurrent states, while the flow backbone enables coherent decoding.
  • Counterfactual Generation. As discussed in the Introduction, DoFlow is among the early approaches for generative counterfactual time-series on a causal DAG, complementing work on treatment effects, post-hoc counterfactual explanations, and causal discovery. Such counterfactual generation is crucial for modeling component relationships and enabling post-hoc analyses in domains like healthcare and hydropower.
6 Conclusions and Discussions
We introduced DoFlow, a flow-based generative framework for causal time-series forecasting that unifies observational, interventional, and counterfactual queries on DAG-structured systems. DoFlow achieves accurate forecasting across causal queries and offers explicit likelihoods that support anomaly detection. Moreover, we provide a supporting counterfactual recovery result (Corollary 4.5) under certain assumptions. Experiments on synthetic DAGs and real applications show strong observational forecasts and effective causal query generation. This framework lays the foundation for domains such as healthcare, energy, and other areas where counterfactual and interventional forecasting is critical. DoFlow can also lead to a broader line of research on flow-based causal generative modeling. Extending this framework to partially observed, multi-modal, or physics-informed systems could enable scientific digital twins that reason under interventions and uncertainty. Such integration of causal structure and generative modeling represents a step toward a general theory of inference and control in complex dynamical environments.
Appendix A Proofs
Proposition 3.1 Given base samples $z_{\tau+1:T}\sim q(\cdot)$, the log-density of the generated time series obtained via the continuous normalizing flow is:
$\log p_{\theta,X_{\tau+1:T}}\!\left(\hat{x}_{\tau+1:T}\mid\hat{H}_{\tau},z_{\tau+1:T}\right)=\sum_{t=\tau+1}^{T}\left[\log q\!\left(z_{t}\right)+\int_{0}^{1}\nabla\!\cdot\,v_{\theta}\bigl(x_{t}(s),s;\,\hat{H}_{t-1}\bigr)\,ds\right].$
Proof. Samples from Continuous Normalizing Flows (CNFs) evolve according to the following Neural ODE:
$\frac{dx_{t}(s)}{ds}=v(x_{t}(s),s;\hat{H}_{t-1}),\quad s\in[0,1],$ (15)
which induces a corresponding evolution of the sample density governed by the Liouville continuity equation:
$\partial p_{\theta}(x_{t},s)+\nabla\cdot\big(p_{\theta}(x_{t},s)v(x_{t},s;\hat{H}_{t-1})\big)=0.$ (16)
Here, $p_{\theta}(x_{t}(s),s)$ denotes the time-dependent probability density of the sample $x_{t}$ at time $t$.
Next, we have that the dynamics of the density $p_{\theta}(\cdot)$ governed by the velocity field $v_{\theta}(x_{t},s;\hat{H}_{t-1})$ is given by:
$\frac{d}{ds}\log p_{\theta}(x_{t}(s),s)$ $=\frac{\nabla p_{\theta}(x_{t}(s),s)\cdot\partial_{s}x_{t}(s)+\partial_{s}p_{\theta}(x_{t},s)}{p_{\theta}(x_{t}(s),s)}$ (17)
$=\frac{\nabla p_{\theta}\cdot\partial_{s}x_{t}(s)+\partial_{s}p_{\theta}}{p_{\theta}}\Big|_{(x_{t}(s),s)}$ (18)
[…]
$=\frac{\nabla p_{\theta}\cdot v-(\nabla p_{\theta}\cdot v+p_{\theta}\nabla\cdot v)}{p_{\theta}}\Big|_{(x_{t}(s),s)}$ (20)
$=-\nabla\cdot v.$ (21)
Starting from an initial sample $z_{t}\sim q(\cdot)$ and integrating from $s=1$ to $s=0$, we have:
$\log p_{\theta,X_{t}}\bigl(x_{t}\mid\hat{H}_{t-1}\bigr)=\log q(z_{t})\;+\;\int_{0}^{1}\nabla_{x}\!\cdot\,v\bigl(x_{t}(s),s;\hat{H}_{t-1}\bigr)\,ds.$ (22)
By summing up all the log-densities within the forecasting window, we obtain:
$\log p_{\theta,X_{\tau+1:T}}\!\left(x_{\tau+1:T}\mid\hat{H}_{\tau},z_{\tau+1:T}\right)=\sum_{t=\tau+1}^{T}\left[\log q\!\left(z_{t}\right)+\int_{0}^{1}\nabla\!\cdot\,v_{\theta}\bigl(x_{t}(s),s;\,\hat{H}_{t-1}\bigr)\,ds\right].$ (23)
Proposition 13 (Encoded as a function of the exogenous noise $U$) Let Assumption 4.1 hold. Without loss of generality, suppose the exogenous noise $U_{t}\sim\mathrm{Unif}[0,1]$. At each time $t$, the observed variable is generated by the structural causal model $X_{t}=f(X_{t-},X_{\mathrm{pa},t-},U_{t})$, and that the flow encoder produces $Z_{t}=\Phi_{\theta}(X_{t};H_{t-1})$. Then there exists a continuously differentiable bijection $g:\mathcal{U}\to\mathcal{Z}$, functionally invariant to $(X_{t-},X_{\mathrm{pa},t-})$, such that,
$Z_{t}\;=\;\Phi_{\theta}\bigl(f(X_{t-},X_{\mathrm{pa},t-},U_{t});H_{t-1}\bigr)\;=\;g\bigl(U_{t}\bigr)\quad\text{a.s.}$ (24)
Proof. Fix a node $i$ (index suppressed) and a time $t$. We write the extended parental state as $S_{t}:=(X_{t-},X_{\text{pa},t-})$. Since the hidden state depends solely on $S_{t}$, we define
$q_{S_{t}}(U_{t})\;:=Z_{t}=\;\Phi_{\theta}\bigl(f(X_{t-},X_{\text{pa},t-},U_{t});\,H_{t-1}\bigr).$
Therefore, our goal becomes proving that $Z_{t}=q_{S_{t}}(U_{t})$ is a function invariant of $S_{t}$. By (A3) in Assumption 4.1, we have $Z_{t}=q_{S_{t}}(U_{t})\perp\!\!\!\perp S_{t}$, and thus:
$p_{Z_{t}|S_{t}=s_{t}}(z_{t})=p_{Z_{t}}(z_{t}).$ (25)
Because continuous normalizing flows are invertible, the encoding function $\Phi_{\theta}:\mathcal{X}\times\mathcal{H}\to\mathcal{Z}$ is invertible with respect to $\mathcal{X}$. In 1-D, this implies monotonicity of $\Phi_{\theta}$ in $X_{t}=f(X_{t-},X_{\text{pa},t-},U_{t})$. Without loss of generality, we assume that $\Phi_{\theta}$ is strictly increasing in $X_{t}=f$. Moreover, by (A2), since $f(\cdot,U_{t})$ is strictly increasing in $U_{t}$, it follows by the composition rule that $q_{S_{t}}(U_{t})$ is strictly increasing in $U_{t}$ and hence bijective in $[0,1]$.
Since $Z_{t}=q_{S_{t}}(U_{t})$, we may apply change of variables formula:
$p_{Z_{t}|S_{t}=s_{t}}(z_{t})\;=\;p_{U_{t}}\!\bigl(q_{s_{t}}^{-1}(z_{t})\bigr)\;\left|\frac{d}{dz_{t}}q_{s_{t}}^{-1}\!\bigl(z_{t}\bigr)\right|=1\cdot\frac{d}{dz_{t}}q_{s_{t}}^{-1}\!\bigl(z_{t}\bigr),$ (26)
where the last equation follows from the uniform distribution of $U_{t}$ and the fact that $p_{Z_{t}}>0$. The absolute value is dropped because of the (WLOG) assumption that $q_{s_{t}}$ is strictly increasing.
Then because of (25), we have that for a fixed $z_{t}$, $\frac{d}{dz_{t}}q_{s_{t}}^{-1}\!\bigl(z_{t}\bigr)$ is the same for any pair $s_{t}=(x_{t-},x_{\text{pa},t-})$. It follows that:
$q_{s_{t}}^{-1}(z_{t})=\int^{z_{t}}\frac{d}{dx}q_{s_{t}}^{-1}\!\bigl(x\bigr)dx=\int^{z_{t}}c(x)dx+c_{s_{t}},$ (27)
where $c(z_{t})=\frac{d}{dz_{t}}q_{s_{t}}^{-1}\!\bigl(z_{t}\bigr)$ is independent of $s_{t}$, and $c_{s_{t}}$ is a constant for each $s_{t}$.
By re-inverting (27), we have:
$q_{s_{t}}(u_{t})=(q_{s_{t}}^{-1})^{-1}(u_{t})=\Bigl(\underbrace{\int^{\bullet}c(x)\,\mathrm{d}x}_{:=G(\bullet)}+c_{s_{t}}\Bigr)^{-1}(u_{t})=G^{-1}\bigl(u_{t}-c_{s_{t}}\bigr).$ (28)
Since $q_{S_{t}}$ is a bijection $0,1]\to\text{supp}(Z_{t})=\{z\in\mathbb{R}:p_{Z_{t}}(z_{t})>0\}$, we have:
$q_{s_{t}}(0)=\inf\ \text{supp}(Z_{t}),\ \ q_{s_{t}}(1)=\sup\ \text{supp}(Z_{t}),\ \text{for any state}\ s_{t}.$ (29)
Therefore,
$q_{s_{t}}(0)=G^{-1}(-c_{s_{t}})=\inf\ \text{supp}(Z_{t}).$ (30)
The support of $Z_{t}$ does not depend on $S_{t}$ because of (25). As a result, for any state $s_{t}=(x_{t-},x_{\text{pa},t-})$, we have that $c_{s_{t}}:=c$ is a constant that does not depend on $s_{t}$.
Therefore, we can write
$g(U_{t})\;:=\;G^{-1}\bigl(U_{t}-c\bigr)\;=\;q_{s_{t}}(U_{t}),\ \forall s_{t}.$ (31)
As a result, we conclude that
$Z_{t}\;=\;\Phi_{\theta}\bigl(X_{t};H_{t-1}\bigr)\;=\;q_{S_{t}}(U_{t})\;=\;g\bigl(U_{t}\bigr)\quad\text{a.s.}.$ (32)
Corollary 4.5 (Counterfactual recovery) Let Assumption 4.1 hold. Consider a factual sample generated by the structural causal model $X_{t}^{\mathrm{F}}=f(X_{t-},X_{\mathrm{pa},t-},U_{t})$, and let its encoded latent be $Z_{t}^{\mathrm{F}}:=\Phi_{\theta}\bigl(X_{t}^{\mathrm{F}};H_{t-1}^{\mathrm{F}}\bigr)$. At time step $t$, we apply the intervention $\mathrm{do}\bigl(X_{t-}=\hat{X}_{t-}^{\mathrm{CF}},X_{\mathrm{pa},t-}=\hat{X}_{\mathrm{pa},t-}^{\mathrm{CF}}\bigr)$, yielding the counterfactual hidden state $\hat{H}_{t-1}^{\mathrm{CF}}$. Then the decoder recovers the true counterfactual at time step $t$ almost surely:
$\hat{X}_{t}^{\mathrm{CF}}:=\Phi_{\theta}^{-1}\!\bigl(Z_{t}^{\mathrm{F}};\,\hat{H}_{t-1}^{\mathrm{CF}}\bigr)=X_{t}^{\mathrm{CF}}.$
Proof. We establish the result by induction over time steps $t$. Following the definition of the true counterfactual value $X_{t}^{\mathrm{CF}}$ defined in (14), we denote the corresponding true hidden state $H_{t-1}^{\mathrm{CF}}$, which is updated using the true values $X_{t-1}^{\mathrm{CF}}$ and $X_{\mathrm{pa},t-1}^{\mathrm{CF}}$.
Base case. Let $t_{0}\in\mathcal{I}$ be the first step after the context window. By the intervention rule,
$X_{t_{0}}^{\mathrm{CF}}=\gamma_{t_{0}}\quad\text{and}\quad\hat{X}_{t_{0}}^{\mathrm{CF}}=\gamma_{t_{0}},$
hence $\hat{X}_{t_{0}}^{\mathrm{CF}}=X_{t_{0}}^{\mathrm{CF}}$. Moreover, since $t_{0}$ is the first intervention, the histories up to $t_{0}-1$ are factual in both constructions, so
$\hat{H}_{t_{0}-1}^{\mathrm{CF}}=H_{t_{0}-1}^{\mathrm{CF}}.$
Induction step. Suppose that for time step until t1t-1, the estimated counterfactual history matches the true one, i.e.,
$\hat{X}_{t-}^{\mathrm{CF}}=X_{t-}^{\mathrm{CF}}.$
Since the counterfactual hidden state is a function of past history until t1t-1, this implies
$\hat{H}_{t-1}^{\mathrm{CF}}=H_{t-1}^{\mathrm{CF}},$ (33)
where the left-hand side is updated recursively from the estimated counterfactual history, and the right-hand side denotes the true hidden state under the true counterfactuals.
By the deterministic and invertible property of the flow for fixed conditioning input, we have
$\Phi_{\theta}^{-1}\!\bigl(\Phi_{\theta}(X_{t};H_{t-1});\,H_{t-1}\bigr)=X_{t}.$ (34)
From the structural causal model, the factual sample is
$X_{t}^{\mathrm{F}}=f(X_{t-},X_{\mathrm{pa},t-},U_{t}),$
for some exogenous noise $U_{t}$. Fixing the same $U_{t}$, and under the intervention $\mathrm{do}(X_{t-}=X_{t-}^{\mathrm{CF}},\,X_{\mathrm{pa},t-}=X_{\mathrm{pa},t-}^{\mathrm{CF}})$, the true counterfactual sample is
$X_{t}^{\mathrm{CF}}=f(X_{t-}^{\mathrm{CF}},X_{\mathrm{pa},t-}^{\mathrm{CF}},U_{t}).$ (35)
By Proposition 4.3, the latent factual representation satisfies
$Z_{t}^{\mathrm{F}}=\Phi_{\theta}\bigl(X_{t}^{\mathrm{F}};\,H_{t-1}^{\mathrm{F}}\bigr)=\Phi_{\theta}\bigl(f(X_{t-},X_{\mathrm{pa},t-},U_{t});\,H_{t-1}^{\mathrm{F}}\bigr)=g(U_{t}),$
which depends only on $U_{t}$ and is invariant to $(X_{t-},X_{\mathrm{pa},t-})$ and thus to $H_{t-1}$. Consequently, under the same intervention $\mathrm{do}(X_{t-}=X_{t-}^{\mathrm{CF}},\,X_{\mathrm{pa},t-}=X_{\mathrm{pa},t-}^{\mathrm{CF}})$, we also have
$\small\Phi_{\theta}\bigl(X_{t}^{\mathrm{F}};\,H_{t-1}^{\mathrm{F}}\bigr)=\Phi_{\theta}\bigr(f(X_{t-},X_{\text{pa},t-},U_{t});H_{t-1}^{\mathrm{F}}\bigr)=g\bigl(U_{t}\bigr)=\Phi_{\theta}\bigr(f(X_{t-}^{\text{CF}},X_{\text{pa},t-}^{\text{CF}},U_{t});H_{t-1}^{\text{CF}}\bigr)=\Phi_{\theta}(X_{t}^{\mathrm{CF}};H_{t-1}^{\text{CF}}),$ (36)
and more simply, we have:
$\Phi_{\theta}\bigl(X_{t}^{\mathrm{F}};\,H_{t-1}^{\mathrm{F}}\bigr)=\Phi_{\theta}(X_{t}^{\mathrm{CF}};H_{t-1}^{\text{CF}}).$ (37)
Combining (33) and (37), we obtain
$\Phi_{\theta}^{-1}\!\bigl(\Phi_{\theta}(X_{t}^{\mathrm{F}};H_{t-1}^{\mathrm{F}});\,\hat{H}_{t-1}^{\mathrm{CF}}\bigr)=\Phi_{\theta}^{-1}\!\bigl(\Phi_{\theta}(X_{t}^{\mathrm{CF}};H_{t-1}^{\mathrm{CF}});\,H_{t-1}^{\mathrm{CF}}\bigr).$ (38)
By (34), the right-hand side of (38) equals $X_{t}^{\mathrm{CF}}$. Since the left-hand side of (38) is precisely the algorithm’s counterfactual encoder-decoder procedure, it follows that
$\hat{X}_{t}^{\mathrm{CF}}:=\Phi_{\theta}^{-1}\!\bigl(Z_{t}^{\mathrm{F}};\,\hat{H}_{t-1}^{\mathrm{CF}}\bigr)=X_{t}^{\mathrm{CF}}.$
Therefore, the decoder recovers the true counterfactual at time step $t$. By induction, the claim holds for all $t$. ∎
Appendix B Preliminaries on Continuous Normalizing Flows
For better logical flow in the main text, we omit the CNF preliminaries there and include them here instead. This section introduces the fundamentals of CNFs and their flow-matching training on general data types.
B.1 Neural ODE and Continuous Normalizing Flow:
A Neural ODE models the evolution of a sample as the solution to an ordinary differential equation (ODE). Concretely, in $\mathbb{R}^{d}$, given an initial condition $x_{0}=x(0)$ at $s=0$, the transformation to the output $x_{1}=x(1)$ at $s=1$ is governed by:
$\frac{dx(s)}{ds}=v(x(s),s),\quad s\in[0,1],$ (39)
where $v:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ is the velocity field parameterized by a neural network. The time horizon is rescaled to $s\in[0,1]$ without loss of generality.
Continuous Normalizing Flow (CNF) is a class of normalizing flows in which the transformation of a probability density is governed by a time-continuous Neural ODE. Let $p(x,s)$ denote the marginal density of $x(s)$. Then $p(x,s)$ evolves according to the Liouville continuity equation implied by (39):
$\partial_{s}p(x,s)+\nabla\cdot\big(p(x,s)\,v(x,s)\big)=0,\quad s\in[0,1],$ (40)
where $\nabla\cdot$ denotes the divergence operator.
When the Neural ODE is well-posed, it induces a continuous and invertible map from the initial sample $x_{0}$ to the terminal sample $x_{1}$. The inverse map is obtained by integrating (39) backward in time. This mechanism allows one to choose $x_{0}\sim p(\cdot,0)$ as the data distribution and $x_{1}\sim p(\cdot,1)$ as a simple base (noise) distribution, typically $N(0,I)$. For convenience, we write $q(\cdot):=p(\cdot,1)$ for the base distribution. Throughout, we use the common choice $q(\cdot)=N(0,I)$.
B.2 Flow Matching
Flow Matching (FM) [41] trains continuous normalizing flows without simulating trajectories by regressing the model velocity $v(x(s),s)$ toward a prescribed target field $u(x(s),s)$. The FM objective is
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{s\sim\mathcal{U}[0,1],\,x\sim p(\cdot,s)}\left[\|v(x(s),s)-u(x(s),s)\|^{2}\right],$ (41)
where $u(x(s),s)$ is analytically specified.
Linear Interpolant: Because $u(x,s)$ is generally intractable unless we condition on the starting point $x_{0}$, Conditional Flow Matching (CFM)  [41] was proposed to find a tractable velocity. Given an observed data $x_{0}$ and an endpoint $x_{1}\sim q(\cdot)$ from base distribution, we can choose an analytic interpolation between $x_{0}$ and $x_{1}$, and define a reference path. An information-preserving and simple choice is the linear interpolant [41]:
$\phi(x_{0},x_{1};s)=(1-s)x_{0}+(s+\sigma_{\text{min}}(1-s))x_{1},$ (42)
where $x_{1}\sim N(0,I)$ and $\sigma_{\text{min}}$ is a small positive hyperparameter ensuring $p(\phi,0)\sim N(x_{0},\sigma_{\text{min}}^{2}I)$. Setting $\sigma_{\text{min}}=0$ recovers the strict linear path $\phi=(1-s)x_{0}+sx_{1}$.
We treat $\phi$ as a fixed reference trajectory that the learned flow is trained to track. Under (42), the associated reference velocity is
$\frac{d\phi}{ds}=(1-\sigma_{\text{min}})x_{1}-x_{0}.$ (43)
Training loss: CFM trains the flow $v(x(s),s)$ by directly regressing it onto the reference velocity field (43). The training objective is given by:
$\mathcal{L}_{\text{CFM}}=\ \mathbb{E}_{s\sim\mathcal{U}[0,1],x_{0}\sim p(\cdot,0),x_{1}\sim q(\cdot)}\|v(\phi,s)-\frac{d\phi}{ds}\|^{2}.$ (44)
Appendix C Data Synthesis
We define four types of causal DAG structures: Tree, Diamond, Fully Connected Layer (FC-Layer), and Chain with skip connections (Chain). For each structure, we design both additive models and nonlinear, non-additive structural causal models.
C.1 Tree
Figure 4: Tree graph over 8 nodes. Exogenous variables $U_{i,t}$ are omitted for clarity but exist for every node at each time $t$. Left: Full node-level causal structure between consecutive time, with all variables $\{X_{1,t},\dots,X_{8,t}\}$ present at each step. Right: Rolled-up (time-suppressed) view over different nodes $\{X_{1},\dots,X_{8}\}$. Each arrow $X_{i}\to X_{j}$ (with $i\neq j$) denotes a lag-1 temporal dependency $X_{i,t-1}\to X_{j,t}$ that holds for all $t$. Both panels depict the same underlying structure.
We consider both additive and nonlinear, non-additive structural causal models:
  • Additive model:
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}+U_{i,t}/4.$ (45)
  • Nonlinear, non-additive model:
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}\,(|U_{i,t}|+0.5)+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}.$ (46)
The root node $X_{1}$ is initialized over the interval $0,t_{0}]$ using a Chain process:
$X_{1,t}=\beta_{1}X_{1,t-1}+A\sin\!\left(\tfrac{2\pi t}{P}+\phi\right)+U_{1,t}.$ (47)
The coefficients $\{\beta_{i}\}$ and $\{\tilde{\beta}_{i}\}$ are selected from predefined sets to ensure that the generated time series contain sufficient noise but do not explode over time. Each parameter is drawn uniformly from its set for every experimental trial, and results are reported as averages across trials.
C.2 Diamond
Figure 5: Diamond graph over 10 nodes. Exogenous variables $U_{i,t}$ are omitted for clarity but exist for every node at each time $t$. Left: Full node-level causal structure between consecutive time, with all variables $\{X_{1,t},\dots,X_{10,t}\}$ present at each step. Right: Rolled-up (time-suppressed) view over different nodes $\{X_{1},\dots,X_{10}\}$. Each arrow $X_{i}\to X_{j}$ (with $i\neq j$) denotes a lag-1 temporal dependency $X_{i,t-1}\to X_{j,t}$ that holds for all $t$. Both panels depict the same underlying structure.
We consider both additive and nonlinear, non-additive structural causal models:
  • Additive model:
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}+U_{i,t}.$ (48)
  • Nonlinear, non-additive model:
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\exp(\beta_{i}X_{i,t-1})\cdot\frac{1}{2+|U_{i,t}|}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}.$ (49)
C.3 Fully Connected Layer (FC-Layer)
Figure 6: FC-Layer graph over 10 nodes. Exogenous variables $U_{i,t}$ are omitted for clarity but exist for every node at each time $t$. Left: Full node-level causal structure between consecutive time, with all variables $\{X_{1,t},\dots,X_{10,t}\}$ present at each step. Right: Rolled-up (time-suppressed) view over different nodes $\{X_{1},\dots,X_{10}\}$. Each arrow $X_{i}\to X_{j}$ (with $i\neq j$) denotes a lag-1 temporal dependency $X_{i,t-1}\to X_{j,t}$ that holds for all $t$. Both panels depict the same underlying structure.
We consider both additive and nonlinear, non-additive structural causal models:
  • Additive
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}+U_{i,t}.$ (50)
  • Nonlinear, non-additive model:
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\sqrt{\frac{1}{2}\left|\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}\right|+|U_{i,t}|}+\beta_{i}X_{i,t-1}.$ (51)
C.4 Chain Linear (Chain)
Figure 7: Chain graph over 50 nodes. Exogenous variables $U_{i,t}$ are omitted for clarity but exist for every node at each time $t$. Left: Full node-level causal structure between consecutive time, with all variables $\{X_{1,t},\dots,X_{50,t}\}$ present at each step. Right: Rolled-up (time-suppressed) view over different nodes $\{X_{1},\dots,X_{50}\}$. Each arrow $X_{i}\to X_{j}$ (with $i\neq j$) denotes a lag-1 temporal dependency $X_{i,t-1}\to X_{j,t}$ that holds for all $t$. Both panels depict the same underlying structure.
We consider both additive and nonlinear, non-additive structural causal models:
  • Additive model:
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}+U_{i,t}.$ (52)
  • Nonlinear, non-additive model:
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\sqrt{\frac{1}{2}\left|\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}\right|+|U_{i,t}|}+\beta_{i}X_{i,t-1}.$ (53)
C.5 Interventional and Counterfactual Simulations
Each time series window has length $T=120$. We set the context window to $\tau=90$ and the forecasting window to $T-\tau=30$. Interventions begin immediately after the context window $\{1,2,\dots,\tau\}$. To obtain system-level interventions, we manipulate the root node(s) $X_{1,t}$, (or $X_{1,t},X_{2,t},X_{3,t}$ in the FC-Layer graph), throughout the forecasting window $\{\tau+1,\dots,T\}$.
To ensure that the intervention produces a substantial deviation from the original observed future, we construct interventions by shifting the root node(s) from earlier periods into the forecasting window, thereby imposing patterns that differ markedly from the original evolution.
Since the underlying structural causal models are known, interventional simulation proceeds by first sampling the exogenous noises $\{U_{i,\tau+1,T}\}_{i=2}^{K}$. The intervened values are then obtained by applying $\mathrm{do}(X_{1,\tau+1:T}=\gamma_{1,\tau+1:T})$ and propagating forward with the sampled noises to generate the intervened trajectories $\{\tilde{X}_{i,\tau+1:T}\}_{i=2}^{K}$. For the FC-Layer graph, the intervened parental set includes $\{X_{i,t}\}_{i=1,2,3}$.
Counterfactual simulation begins by recovering the exogenous noises $\{U_{i,t}\}_{i>1,\,t\in[\tau+1,T]}$ from the observed factual future $\{X_{i,\tau+1:T}^{\mathrm{F}}\}_{i=2}^{K}$ using the structural causal models. Next, the intervention $\mathrm{do}(X_{1,\tau+1,T}=\gamma_{1,\tau+1:T})$ is applied, and the system is propagated forward with the recovered noises to generate the counterfactual trajectories $\{X_{i,\tau+1:T}^{\mathrm{CF}}\}_{i=2}^{K}$. For the FC-Layer graph, the intervened parental set includes $\{X_{i,t}\}_{i=1,2,3}$.
C.6 Metrics
We evaluate model performance using both Root Mean Squared Error (RMSE) and Maximum Mean Discrepancy (MMD). Let the test batch size be $B=128$, the groundt-ruth value for node $i$ at time $t$ in batch $b$ be $x_{i,t}^{(b)}$, and the corresponding model prediction be $\hat{x}_{i,t}^{(b)}$.
For each test sequence $b$, given the same context $\{x_{i,1:\tau}^{(b)}\}_{i=1}^{K}$, we generate $N=100$ realizations of both the model-estimated and the true observational/interventional forecasting trajectories, in order to obtain more accurate evaluation metrics.
To ensure comparability across different scales of time series, we apply standard normal scaling to each batch $b$ over the forecasting window, using the mean $\mu_{i}^{(b)}$ and standard deviation $\sigma_{i}^{(b)}$ computed from its context window of node $i$. The generated time series dataset has a total length of 20,000, with a stride of 1, resulting in 15,881 training samples (80%) and 3,881 testing samples (20%).
The RMSE of node $i$ for a single realization is defined as
$\mathrm{RMSE}_{i}=\sqrt{\frac{1}{(T-\tau)B}\sum_{t=\tau+1}^{T}\sum_{b=1}^{B}\bigl(\hat{x}_{i,t}^{(b)}-x_{i,t}^{(b)}\bigr)^{2}}.$ (54)
We run each algorithm 50 times, each time using different batches with batch size $B=128$. The final reported RMSE is obtained by first averaging over all nodes $i=1,\dots,K$, then averaging across the $N=100$ realizations, and finally averaging over the 50 runs. The reported standard deviation (std) is calculated from the variation across the 50 runs.
To calculate MMD, we first flatten each trajectory (length $T-\tau$, dimension $D$) into a vector in $\mathbb{R}^{(T-\tau)D}$. The sample size of both the true and the estimated trajectories $\{x_{i,\tau+1:T}\}_{i=1}^{K}$ is $BN$, since for each batch $b$ we simulate $N=100$ realizations. The empirical MMD is then defined as
$\widehat{\mathrm{MMD}}^{2}=\frac{1}{(BN)(BN-1)}\!\sum_{\begin{subarray}{c}a,a^{\prime}=1\\ a\neq a^{\prime}\end{subarray}}^{BN}k(x_{a},x_{a^{\prime}})\;+\;\frac{1}{(BN)(BN-1)}\!\sum_{\begin{subarray}{c}b,b^{\prime}=1\\ b\neq b^{\prime}\end{subarray}}^{BN}k(\tilde{x}_{b},\tilde{x}_{b^{\prime}})\;-\;\frac{2}{(BN)^{2}}\!\sum_{a=1}^{BN}\sum_{b=1}^{BN}k(x_{a},\tilde{x}_{b}),$ (55)
where we use the Gaussian kernel
$k(x,x^{\prime})=\exp\!\Big(-\tfrac{\|x-x^{\prime}\|^{2}}{2\sigma^{2}}\Big),$
and the bandwidth $\sigma$ is chosen via the pooled median heuristic.
Appendix D Comparisons to BGM settings
Bijective Generation Mechanisms (BGM). The BGM framework [47] shows that if the true structural mechanism $f$ is in the BGM class (bijective/strictly monotone in the exogenous noise), then any learned mechanism $\hat{f}$ that (i) is also in the BGM class and (ii) matches the observed distribution—i.e., for the same parents value $X$ and action $A$,
$\small\hat{f}(X_{t-},X_{\text{pa},t-},U_{t})$ $\stackrel{{\scriptstyle d}}{{=}}\ f(X_{t-},X_{\text{pa},t-},U_{t})$
$i.e.,(\hat{f}(X_{t-},X_{\text{pa},t-},\cdot))_{\#}P_{U}$ $=(f(X_{t-},X_{\text{pa},t-},\cdot))_{\#}P_{U},$
yields the same counterfactuals as $f$. Their result is model-agnostic, at the class level.
DoFlow. Our findings are specific to a particular methodology and architecture, as this is a methodology paper where we analyze continuous normalizing flows (CNFs). With the additional assumption (A3) specific to DoFlow, Proposition 4.3 shows that the DoFlow’s encoded latent $Z_{t}$ is bijective in the exogenous noise $U_{t}$. Consequently, Corollary 4.5 proves that the encode–decode procedure recovers the true counterfactual.
Relationships. Because of Proposition 4.3 (enabled by the CNF), DoFlow implements a bijective-in-noise mechanism and is thus comparable to BGM in the sense that $\hat{f}$ is bijective in $U_{t}$. However, they are not the same: BGM additionally requires observational distribution matching to obtain class-level identifiability, whereas the proofs of Proposition 4.3 and Corollary 4.5do not assume such matching.
Alternative route. The counterfactual recovery result (Corollary 4.5) can alternatively be established by imposing an additional assumption:
(A4) Observational matching. For each $(X_{t-},X_{\text{pa},t-})$, the DoFlow–induced observational law matches the true one, i.e.
$(\hat{f}(X_{t-},X_{\text{pa},t-},\cdot))_{\#}P_{U}=(f(X_{t-},X_{\text{pa},t-},\cdot))_{\#}P_{U}.$
Under (A1)–(A3), Proposition 4.3 establishes a continuously differentiable bijection $g:U_{t}\to Z_{t}$ (independent of $(X_{t-},X_{\text{pa},t-}$), so the induced mechanism $\hat{f}$ implemented by DoFlow is bijective in the exogenous noise, i.e., DoFlow lies in the BGM class. With the additional assumption (A4), we can therefore directly invoke the counterfactual recovery result of the BGM framework [47].
Appendix E More Experimental Results
E.1 Empirical Validation of Assumption (A3)
We test (A3): $p_{\theta}(Z_{t}\mid H_{t-1})=q(Z_{t})=N(Z_{t};0,1)$, using a two-sample Maximum Mean Discrepancy (MMD). Specifically, we test the joint factorization $p(Z_{t},H_{t-1})=q(Z_{t})p(H_{t-1})$ by computing the MMD between the empirical joint sample $\{(z_{t},h_{t-1})\}$ and a synthetic joint sample $\{(z_{t}^{\prime},h_{t-1})\}$, where $z_{t}^{\prime}\sim q=N(0,1)$.
We employ a product kernel $k((z,h),(z^{\prime},h^{\prime}))=k_{Z}(z,z^{\prime})\cdot k_{H}(h,h^{\prime})$, where both $k_{Z}(z,z^{\prime})$ and $k_{H}(h,h^{\prime})$ are radial basis function (RBF) mixtures:
$k_{Z}(z,z^{\prime})$ $=\exp(-\frac{\|z-z^{\prime}\|^{2}}{2\sigma_{Z}^{2}}),$ (56)
$k_{H}(h,h^{\prime})$ $=\exp(-\frac{\|h-h^{\prime}\|^{2}}{2\sigma_{H}^{2}}).$
To set the bandwidths $\{\sigma_{Z}\}$ and $\{\sigma_{H}\}$, we first compute the pooled pairwise distances on the union of the observed and synthetic sets (e.g., $\tilde{Z}=Z\cup Z^{\prime}$). We then calculate the median of these distances, and define the bandwidths as $\tfrac{1}{2}\,\text{median}$. The empirical MMD is given as:
$\widehat{\mathrm{MMD}}^{2}_{u}=\frac{1}{n(n-1)}\sum_{i\neq j}k(x_{i},x_{j})+\frac{1}{n(n-1)}\sum_{i\neq j}k(y_{i},y_{j})-\frac{2}{n(n-1)}\sum_{i\neq j}k(x_{i},y_{j}),$ (57)
with $x_{i}=(z_{i},h_{i})$, $y_{i}=(z_{i}^{\prime},h_{i})$.
For more reliable comparisons, we additionally sample two groundt-ruth sets, $\{(z_{t}^{\prime},h_{t-1})\}$ and $\{(z_{t}^{\prime\prime},h_{t-1})\}$, where both $z_{t}^{\prime}\sim\mathcal{N}(0,1)$ and $z_{t}^{\prime\prime}\sim\mathcal{N}(0,1)$. The test is conducted on the linear additive “Tree” DAG, performed separately for each dimension, and the aggregated results are reported in the table below.
Table 2: Comparison of DoFlow and groundt-ruth samples under the Tree DAG using per-dimension MMD. Reported are the mean, minimum, and maximum values across dimensions.
Mean Min Max
DoFlow $3.7\times 10^{-2}$ $5.0\times 10^{-3}$ $8.3\times 10^{-2}$
True $2.4\times 10^{-2}$ $4.7\times 10^{-3}$ $6.9\times 10^{-2}$
As shown in Table 2, DoFlow achieves small two-sample joint MMD values, which are comparable to those obtained from the groundt-ruth independent samples $\{(z_{t}^{\prime},h_{t-1})\}$ and $\{(z_{t}^{\prime\prime},h_{t-1})\}$. This provides empirical support for the validity of Assumption (A3).
E.2 Computational Costs
The number of training samples is 15,881 for each simulated datasets. We use a batch size of 128, and all experiments—including training and sampling time comparisons—are conducted on a single A100 GPU.
Table 3: Comparison of model size.
Tree Diamond Layer Chain Hydropower
DoFlow $94,664$ $121,658$ $133,946$ $646,178$ $99,592$
GRU $36,830$ $50,106$ $51,732$ $124,412$ $52,026$
TFT $94,860$ $100,472$ $124,775$ $531,026$ $95,318$
TiDE $108,301$ $118,332$ $118,332$ $544,539$ $122,462$
TSMixer $117,818$ $119,318$ $119,318$ $550,177$ $110,518$
DeepVAR $73,712$ $89,280$ $89,280$ $184,280$ $84,290$
MQF2 $128,677$ $160,846$ $160,846$ $702,681$ $125,459$
Table 4: Comparison of training time per epoch, epochs to convergence, and sampling time (for 1,000 forecast series over a 30-step horizon) across Tree, Diamond, and Layer structures.
Methods Training Time / Epoch Epochs to Conv. Total Training Time Sampling Time
DoFlow $42.3$s $10$ 8.098.09min $9.45$s
GRU $17.8$s $15$ 5.035.03min $6.80$s
TFT $21.0$s $15$ 5.285.28min $11.07$s
TiDE $29.3$s $25$ 12.112.1min $2.23$s
TSMixer $27.9$s $25$ 12.112.1min $8.83$s
DeepVAR $20.6$s 2020 11.6511.65min $7.86$s
MQF2 $49.8$s 2020 16.616.6min $19.61$s
E.3 Synthetic Data Experiments
Table 5: MMD for observational, interventional, and counterfactual time series forecasting across causal structures: Tree, Diamond, FC-Layer, and Chain. Reported values are averaged over 50 test batches, each containing 128 test series.
Tree Diamond FC-Layer Chain
Additive NLNA Additive NLNA Additive NLNA Additive NLNA
Obs. Int. Obs. Int. Obs. Int. Obs. Int. Obs. Int. Obs. Int. Obs. Int. Obs. Int.
[…]
TFT $\mathbf{0.08_{\pm.01}}$ $0.13_{\pm.04}$ $0.18_{\pm.05}$ $25_{\pm.09} 0.0$$08_{\pm.03}$ $0.15_{\pm.05}$ $0.20_{\pm.05}$ $26_{\pm.10} 0.0$$06_{\pm.02}$ $0.18_{\pm.06}$ $0.20_{\pm.06}$ $0.29_{\pm.13}$ $0.13_{\pm.03}$ $0.19_{\pm.09}$ $0.23_{\pm.05}$ $0.30_{\pm.12}$
[…]
DeepVAR $0.12_{\pm.02}$ $0.10_{\pm.02}$ $0.18_{\pm.03}$ $0.19_{\pm.03}$ $0_{\pm.02} 0.11±.0$$1_{\pm.02}$ $0.22_{\pm.04}$ $25_{\pm.08} 0.0$$08_{\pm.02}$ $0.23_{\pm.04}$ $0.26_{\pm.06}$ $0.35_{\pm.10}$ $0.11_{\pm.02}$ $0.16_{\pm.03}$ $0.24_{\pm.05}$ $0.29_{\pm.05}$
MQF2 $\mathbf{0.08_{\pm.01}}$ $0.16_{\pm.03}$ $0.20_{\pm.03}$ $29_{\pm.06} 0.0$$09_{\pm.02}$ $0.15_{\pm.04}$ $0.19_{\pm.03}$ $30_{\pm.10} 0.0$$07_{\pm.01}$ $0.20_{\pm.03}$ $0.21_{\pm.03}$ $0.34_{\pm.09}$ $0.17_{\pm.03}$ $0.22_{\pm.05}$ $0.28_{\pm.04}$ $0.33_{\pm.07}$
Table 6: RMSE for observational, interventional, and counterfactual time-series forecasting on the Chain causal structure.
Chain
Additive NLNA
Obs. Int. CF. Obs. Int. CF.
DoFlow $\mathbf{0.61_{\pm.13}}$ $\mathbf{0.62_{\pm.13}}$ $\mathbf{0.69_{\pm.15}}$ $\mathbf{0.69_{\pm.16}}$ $\mathbf{0.74_{\pm.17}}$ $0.75_{\pm.17}$
GRU $0.68_{\pm.11}$ $1.01_{\pm.14}$ NA $0.80_{\pm.11}$ $1.21_{\pm.16}$ NA
TFT $0.63_{\pm.17}$ $1.10_{\pm.24}$ NA $0.78_{\pm.15}$ $1.29_{\pm.19}$ NA
TiDE $0.65_{\pm.10}$ $1.07_{\pm.17}$ NA $0.77_{\pm.12}$ $1.16_{\pm.18}$ NA
TSMixer $0.67_{\pm.11}$ $1.09_{\pm.18}$ NA $0.75_{\pm.14}$ $1.20_{\pm.20}$ NA
DeepVAR $\mathbf{0.62_{\pm.12}}$ $0.97_{\pm.18}$ NA $0.84_{\pm.13}$ $1.09_{\pm.16}$ NA
MQF2 $0.73_{\pm.13}$ $1.18_{\pm.19}$ NA $0.90_{\pm.15}$ $1.30_{\pm.21}$ NA
Figure 8: Left:Tree” interventional forecasting results. Node $X_{1,t}$ is intervened. DoFlow provides 50% and 90% prediction intervals; the orange lines indicate the true interventional future. Right:Layer” counterfactual forecasting results. Nodes $X_{1,t}$, $X_{2,t}$, and $X_{3,t}$ are intervened. DoFlow provides a single forecast in green; the orange lines indicate the true counterfactual future.
E.4 Hydropower System
Figure 9: Hydropower system graph over 8 nodes. Exogenous variables $U_{i,t}$ are omitted for clarity but exist for every node at each time $t$. Left: Full node-level causal structure between consecutive time, with all variables $\{X_{1,t},\dots,X_{8,t}\}$ present at each step. Right: Rolled-up (time-suppressed) view over different nodes $\{X_{1},\dots,X_{8}\}$. Each arrow $X_{i}\to X_{j}$ (with $i\neq j$) denotes a lag-1 temporal dependency $X_{i,t-1}\to X_{j,t}$ that holds for all $t$. Both panels depict the same underlying structure.
In Figure 11, $X_{1}$ and $X_{2}$ denote the horizontal and vertical rotational vibrations of the hydraulic turbine, which directly drive the generator’s operation. The generator’s horizontal and vertical dynamics are represented by $X_{3}$ and $X_{4}$, while $X_{5}$ captures the generator’s current output delivered to the transformer ($X_{8}$) for voltage regulation and transmission to the power grid. The Metering and Control (M&C) unit ($X_{7}$) continuously monitors the generator’s performance and turbine stability to ensure coordinated operation. In addition, $X_{6}$ represents an auxiliary generator state (thermal signal) recorded independently for diagnostic purposes but not directly connected in the causal structure.
Table 7 reports the RMSE results for observational and interventional time-series forecasting in the hydropower system. For interventional forecasting, we use 12 real power outages where the root nodes Turbine $X_{1}$ and $X_{2}$ fail, causing the entire system to an outage. In this setting, the root nodes are treated as intervened by the breakdown signals. The reported averages and standard deviations are computed over 12 runs with batch size $B=1$ for the interventional case, in contrast to 50 runs with batch size $B=128$ for the observational case.
Notably, the hydropower signals are highly unstable, with turbine flow and generator readings often exhibiting abrupt jumps or burnouts without clear patterns. Consequently, all methods face difficulty in accurate prediction, and the relatively high RMSE values reflect this inherent challenge. Nevertheless, our model performs consistently better than others under these conditions.
Hydropower System
Obs. Int.
DoFlow $\mathbf{1.13_{\pm.18}}$ $\mathbf{1.21_{\pm.19}}$
GRU $2.05_{\pm.32}$ $2.45_{\pm.35}$
TFT $1.82_{\pm.25}$ $2.16_{\pm.41}$
TiDE $1.49_{\pm.24}$ $2.08_{\pm.40}$
TSMixer $1.51_{\pm.25}$ $2.11_{\pm.32}$
DeepVAR $1.78_{\pm.26}$ $2.39_{\pm.28}$
MQF2 $1.97_{\pm.24}$ $2.62_{\pm.34}$
Table 7: RMSE for observational and interventional time-series forecasting in the hydropower system.
Figure 10: Anomaly detection by DoFlow on real power outages in the hydropower system (two segments shown).
Figure 10 presents two segments of DoFlow log-probabilities during outage time windows. In each case, minute 50 marks the actual outage event. Notably, DoFlow’s log-probability output becomes abnormal well before the outage occurs—sometimes as early as 20 minutes prior (e.g., $X_{6,t}$ in the right panel) and as late as 10 minutes prior. This allows anomalies to be detected in advance of the true outage.
E.5 Cancer Treatment Effects
We apply DoFlow to interventional forecasting of cancer tumor outcomes. For each test patient, the model observes the first $55$ days of factual history $\{(\mathbf{X}_{t},Y_{t})\}_{t=1}^{55}$, where $\mathbf{X}_{t}=\{X_{i,t}\}_{i=1}^{4}$ represents chemotherapy and radiotherapy assignments and dosages, and $Y_{t}$ denotes the tumor volume.
During the forecasting window (days 56566262), the treatment schedule is replaced by one of ten pre-defined intervention plans $\mathcal{I}^{j}=\{(i,t):X_{i,t}\leftarrow m_{i,t}^{j}\}$, where $j$ indicates the $j$-th treatment plan. At each time step $t$, DoFlow estimates the interventional tumor outcome by sampling from the learned flow model:
$\hat{Y}_{t}=\Phi^{-1}_{\theta}(z_{t};\hat{H}_{t-1}),\qquad z_{t}\sim\mathcal{N}(0,I),$ (58)
where $\Phi^{-1}_{\theta}$ denotes the learned reverse flow conditioned on the recurrent hidden state $\hat{H}_{t-1}$, which encodes the patient’s historical outcomes and past treatments. The hidden state is updated autoregressively using the newly generated $\hat{Y}_{t}$ and the active treatments $\mathbf{X}_{t}=\mathbf{m_{t}^{j}}$.
Figure 11: Cancer Treatment DAG over 8 nodes. Exogenous variables $U_{i,t}$ are omitted for clarity but exist for every node at each time $t$. Left: Full node-level causal structure between consecutive time, with all treatment variables $\{X_{i,t}\}_{i=1}^{4}$ and cancer tumor outcome $Y_{t}$ present at each step. Right: Rolled-up (time-suppressed) view over different nodes. Each arrow $X_{i}\to Y$ denotes a lag-1 temporal dependency $X_{i,t-1}\to Y_{t}$ that holds for all $t$. Both panels depict the same underlying structure.
Table 8: Normalized RMSEτ for causal treatment effects on cancer tumor outcome. At future step $\tau$, RMSE is computed across all patient–option pairs as. Column groups represent the chemotherapy and radiotherapy application budgets $(\gamma_{c},\gamma_{r})$ in various data simulation scenarios.
  $\gamma_{c}=5,\ \gamma_{r}=5$ $\gamma_{c}=5,\ \gamma_{r}=0$ $\gamma_{c}=0,\ \gamma_{r}=5$
$\tau$ DoFlow CRN RMSN MSM DoFlow CRN RMSN MSM DoFlow CRN RMSN MSM
3 $\mathbf{1.25\%}$ 2.43% 3.16% 6.75% $\mathbf{0.49\%}$ 1.08% 1.35% 3.68% $\mathbf{0.94\%}$ 1.54% 1.59% 3.23%
4 $\mathbf{1.73\%}$ 2.83% 3.95% 7.65% $\mathbf{0.76\%}$ 1.21% 1.81% 3.84% $\mathbf{1.10\%}$ 1.81% 2.25% 3.52%
5 $\mathbf{2.08\%}$ 3.18% 4.37% 7.95% $\mathbf{0.88\%}$ 1.33% 2.13% 3.91% $\mathbf{1.27\%}$ 2.03% 2.71% 3.63%
6 $\mathbf{2.74\%}$ 3.51% 5.61% 8.19% $\mathbf{1.09\%}$ 1.42% 2.41% 3.97% $\mathbf{1.69\%}$ 2.23% 2.73% 3.71%
7 $\mathbf{3.22\%}$ 3.93% 6.21% 8.52% $\mathbf{1.33\%}$ 1.53% 2.43% 4.04% $\mathbf{2.01\%}$ 2.43% 2.88% 3.79%
We compute the normalized root mean-squared error (RMSE) at the $\tau$-th step across all patients and treatment options as
$\text{RMSE}_{\tau}=\frac{\sqrt{\frac{1}{N}\sum_{n=1}^{N}\left(Y_{n,t+\tau}-\hat{Y}_{n,t+\tau}\right)^{2}}}{\frac{1}{N}\sum_{n=1}^{N}Y_{n,t+\tau}},$ (59)
where $N$ is the total number of patient–option pairs in the test set. Smaller NRMSE indicates more accurate estimation of causal treatment effects.
Table 8 reports the normalized RMSE results for causal treatment effect estimation on cancer tumor outcomes. We compare our method with three established baselines: CRN (Counterfactual Recurrent Network) [5], RMSN (Recurrent Marginal Structural Network) [35], and MSM (Marginal Structural Model) [43]. The datasets follow the construction in Bica et al. [5], where $\gamma_{c}$ and $\gamma_{r}$ denote the treatment-application budgets for chemotherapy and radiotherapy, respectively. Some baseline results in Table 8 are adopted directly from Bica et al. [5].